Choose a,b,c,d
4q + 4q2 and SMC = 4 + 8q where q is output and all fixed costs are sunk, the firms short-run supply curve is A) q-4+8p for p24 and zero otherwise. B) q-4+4p for p24 and zero otherwise. C)q-o if poz and q-1/4p-i if pz2 D) q=o if pc4 and qa/8p-1/2 if p24
Answer #1

9. Short Run Total Cost (STC) = 300 + 4q + 4q2

Short Run Marginal Cost (SMC) = 4 + 8q

Lets, first take the derivative of STC =  (STC)/aq = 4 + 8q which is equal to Short Run Marginal Cost (SMC)

again taking the second derivative of STC we get = partial (STC)^{2}/partial q = 4

The firm always makes production decisions based on the Marginal Cost curve. So, the firm would supply only when P>= 4 and would not supply anything below that price. Hence, the firm's short run supply curve is Option A) q = 4 + 8p for p>=4 and zero otherwise.